### v1

@@ Abstract
We report on a direct search for elastic photon-photon scattering using x-ray and $\gamma$ photons from a laser-plasma based experiment. A gamma photon beam produced by a laser wakefield accelerator provided a broadband gamma spectrum extending to above $E_{\gamma}=200$ MeV. These were collided with a dense x-ray field produced by the emission from a laser heated germanium foil at $E_{x}\approx 1.4$ keV, corresponding to an invariant mass of $\sqrt{s}=1.22\pm 0.22$ MeV. In these asymmetric collisions elastic scattering removes one x-ray and one high-energy $\gamma$ photon and outputs two lower energy $\gamma$ photons. No changes in the $\gamma$ photon spectrum were observed as a result of the collisions allowing us to place a 95% upper bound on the cross section of $1.5\times 10^{15}$ µb. Although far from the QED prediction, this represents the lowest upper limit obtained so far for $\sqrt{s}\lesssim 1$ MeV.
@@ 1 Introduction
This paper reports on a search for elastic photon-photon scattering at $\sqrt{s}\approx 1$ MeV. The search is motivated at this energy scale because it is relevant to astrophysics, it is where the cross-section takes its maximum value, and at this energy, the QED effect has only very weak experimental bounds. This scale also includes energies over the threshold for creation of real electron-positron pairs via the linear Breit-Wheeler process, which therefore is a sub-process of photon-photon scattering at these energies (this can be understood by the optical theorem [29]). Linear Breit-Wheeler is also being searched for at PW-class optical lasers [30, 31], and multi-photon Breit-Wheeler pair-creation forms part of the science goals for the LUXE experiment planned at DESY [32] and the E320 experiment at SLAC [33].
@@
In this asymmetric collision, a scattering event typically removes one x-ray and one high energy $\gamma$ photon from the beam and replaces it with two lower energy $\gamma$ photons. If sufficient numbers of $\gamma$ photons were to scatter in the x-ray field, we would therefore be able to detect the effect in the photon energy spectrum. This is illustrated in Fig. 1 for the idealised case of a narrow energy spread 500 MeV $\gamma$ photon beam scattering from a dense x-ray field. If however, insufficient scattering events occur to produce a measurable change in the spectrum, this allows us to place a limit on the photon-photon scattering cross-section.
@@
Figure 1: Illustration of the effect on the photon energy spectrum of copious photon-photon scattering of a 500 MeV $\gamma$ photon beam in a dense 1.5 keV x-ray field.
@@ 2 Experimental set-up
One of the Gemini laser pulses was used to generate a dense x-ray field by rapidly heating a 100 nm germanium (Ge) foil. As this solid Ge foil is heated, it turns into a plasma, leading to the emission of intense x-ray radiation predominantly due to M-L band transitions [52, 53]. The heating laser pulse had a duration of 40 ps (fwhm intensity) and a total energy of $10.7\pm 0.3$ J. It was focused to an elliptical spot using a distributive phase plate, with major and minor axes of $(217\pm 6)$ µm and $(77\pm 6)$ µm respectively, which contained 72% of the total energy. The Ge targets were mounted on a Kapton (C22H10N2O5) tape with a lower average atomic number, limiting the mass of Ge close to the interaction which is a potential noise source. A motorised tape-drive was used to change targets between shots.
To diagnose the x-ray field, a pinhole imaging system and crystal spectrometer were used. The pinhole imaging system gave an on-shot measure of both the emission spot size and the target alignment. The spectrometer used a flat, thallium acid phthalate (TlAP) crystal, with a spectral window of $\approx 700$ eV, centred at approximately $1.6$ keV, signal above approximately $1.5$ keV was surpressed due to an aluminium filter. This spectral window is around the M-L band transitions of Ge.
@@ 2.2 $\gamma$ photon source
The $\gamma$ photon source was generated through bremsstrahlung emission, which first requires a beam of high energy electrons. To produce these electrons, one of the Gemini laser beams was focused into a 17.5 mm gas cell filled with helium and a 2% nitrogen dopant. The duration of the laser pulse was $45\pm 5$ fs (fwhm intensity) and the focal spot was $(44\pm 2)$ µm $\times$ $(53\pm 2)$ µm (fwhm intensity). The laser energy on target was $5.5\pm 0.6$ J, corresponding to a normalised vector potential $a_{0}=1.1\pm 0.2$. Through the laser wakefield acceleration mechanism [34], a beam of high energy electrons (energy up to $\approx 800$ MeV, charge $\approx 50$ pC ) was emitted from the gas cell. These electrons then passed through a 0.5 mm thick bismuth (Bi) foil, acting as a bremsstrahlung converter. This emits a beam of high energy $\gamma$ photons with a similar duration to that of the driving laser pulse.
@@
Bremsstrahlung converters also generate a large number of low energy, divergent $\gamma$ photons, which are a potential noise source. If these divergent $\gamma$ photons were to interact with the Ge foil, or another part of the experimental setup, they would produce background through the Compton scattering process. To prevent this a 100 mm block of tungsten (W) with a 2 mm diameter hole drilled through the centre was used to collimate the $\gamma$ photon beam. This collimator effectively removes any $\gamma$ photon travelling at an angle greater than 10 mrad to the beam propagation direction.
@@
To reduce the background further, a 50 mm tungsten block was placed just off-axis, shadowing the Ge foil from the $\gamma$ photon beam. Placing such a high Z-material close the $\gamma$ photon beam axis will itself generate a large number of background Bethe–Heitler pairs which would pass through the x-ray field and create a background signal through Compton scattering. Therefore, a 30 cm dipole magnet with a field strength of $B=1$ T, was used to remove these before the photon-photon interaction zone. A CsI crystal screen was placed before the $\gamma$ spectrometer to detect the footprint of the $\gamma$ beam, but is not used for this analysis.
@@ 2.3 $\gamma$ photon spectrometer
The $\gamma$ photon spectrometer consisted of a stack of $33\times 47$ CsI crystals doped with thallium and imaged onto a 14-bit EMCCD camera. This was positioned inside the $\gamma$ photon beam path, with higher energy $\gamma$ photons penetrating deeper into the stack than lower energy $\gamma$ photons. The $\gamma$-photon spectrum is determined using a trial function based forward model. Bayesian inference is used to determine the best fit parameters of the trial function and their uncertainty.
The detector was calibrated to remove systematic effects such as variations in light yield in each crystal due to crystal imperfections and misalignment or effects produced by the imaging system. We compared the measured energy deposited in each crystal averaged over a large number of shots with that predicted in Geant4 using the average electron energy spectrum (as measured on a series of shots without the bismuth foil intercepting the electron beam) and generated a correction factor which can then be applied to each column of crystals.
@@
With the likelihood and priors set, we can use equation 3 to calculate the posterior. Given that the numerator involves a three-dimensional integral, it is most efficiently solved using a Markov chain Monte Carlo (MCMC) method. In figure 3 we can see an example of this calculation performed on a randomly selected shot from the data set.
Figure 3: a) Gamma photon spectrum (Geant4 simulations based on measured electron beam) b) Measured x-ray photon spectrum c) Measured signal on the CsI array from a randomly selected null shot (red circles). The CsI signal generated by the forward model of the detector for the average (black) and 100 randomly selected samples (grey) samples from the posterior distribution of $E_{c}$. The mean value corresponds to $E_{c}=83.2\pm 3.2$ MeV.
@@ 3.1 $\gamma$ photon spectrum
We compare the distribution of $E_{c}$ on collision and null shots in various ways. Figure 4 shows the histogram of the inferred value for the $\gamma$ photon spectrum $E_{c}$. The relatively small number of shots in each distribution means it is not immediately clear if differences between them are significant.
Figure 4: Distribution of inferred value for the $\gamma$ photon spectrum $E_{c}$ for 32 null shots (blue) and 22 collision shots (red)
To investigate this, figure 5 shows the distribution of mean and standard deviation calculated from 100,000 bootstrap samples of the null and collision shot data. The fact that the distributions overlap illustrates further that there is no significant difference between the distribution of $E_{c}$ on null and collision shots.
Figure 5: Bootstrap estimate of (a) the mean and (b) the standard deviation of $E_{c}$ for null shots (blue) and collision shots (red)
Figure 6 shows the empirical cumulative distribution function (ecdf) of $E_{c}$ for the data. Also shown are 50 ecdfs for bootstrap samples of the data, these effectively represent the uncertainty in the ecdfs. The overlap between these ecdfs again illustrates that there is no significant difference between the distributions.
Figure 6: The empirical cumulative distribution (ecdf) function of $E_{c}$ for null shots (blue) and collision shots (red). The thick lines show the ecdf for the experimental data, the thin lines show ecdfs calculated from 50 bootstrap samples of the data.
@@
where $F_{\mathrm{N}}\left(E_{c}\right)$ and $F_{\mathrm{C}}\left(E_{c}\right)$ are the cumulative distribution functions for the null and collision shots respectively. The null hypothesis is accepted at the 95% confidence level if $D<0.378$[54] for a data sets with 32 and 22 samples. The value obtained for our data set is D = 0.216, so we cannot reject the null hypothesis.
We can also calculate the two-sample Kolmogorov-Smirnov test statistic for a large number of bootstrap samples from the data to estimate the uncertainty in the KS test statistic. This is shown in figure 7. It shows that the bulk of the distribution ($\approx 90\%$) lies below the critical value, providing further strong evidence that we cannot reject the null hypothesis, i.e. we must assume that collisions between $\gamma$ photons and the dense x-ray field did not produce a detectable difference in the energy spectrum of the $\gamma$ photons.
Figure 7: Distribution of the two sample Kolmogorov-Smirnov test statistic comparing the distribution of $E_{c}$ for null shots and collision shots from 100,000 bootstrap samples from the data. The red line shows the value above which the null hypothesis (that the null and collision shots are from the same distribution) can be rejected at the $\alpha=0.05$ significance level.
@@ 4 Conclusions
Data will be made available on request.

### v2

@@ Abstract
We report on a direct search for elastic photon-photon scattering using x-ray and $\gamma$ photons from a laser-plasma based experiment. A $\gamma$ photon beam produced by a laser wakefield accelerator provided a broadband $\gamma$ spectrum extending to above $E_{\gamma}=200$ MeV. These were collided with a dense x-ray field produced by the emission from a laser heated germanium foil at $E_{x}\approx 1.4$ keV, corresponding to an invariant mass of $\sqrt{s}=1.22\pm 0.22$ MeV. In these asymmetric collisions elastic scattering removes one x-ray and one high-energy $\gamma$ photon and outputs two lower energy $\gamma$ photons. No changes in the $\gamma$ photon spectrum were observed as a result of the collisions allowing us to place a 95% upper bound on the cross section of $1.5\times 10^{15}$ µb. Although far from the QED prediction, this represents the lowest upper limit obtained so far for $\sqrt{s}\lesssim 1$ MeV.
@@ 1 Introduction
This paper reports on a search for elastic photon-photon scattering at $\sqrt{s}\approx 1$ MeV. This is a $2\to 2$ process with two free photons in the out state, in contrast to the $1\to 1$ stimulated photon scattering process searched for in cavity experiments (see e.g. [29] for a comparison). The search at this energy is motivated at this energy scale because it is relevant to astrophysics, it is where the cross-section takes its maximum value, and at this energy, the QED effect has only very weak experimental bounds. This scale also includes energies over the threshold for creation of real electron-positron pairs via the linear Breit-Wheeler process, which therefore is a sub-process of photon-photon scattering at these energies (this can be understood by the optical theorem [30]). Linear Breit-Wheeler is also being searched for at PW-class optical lasers [31, 32], and multi-photon Breit-Wheeler pair-creation forms part of the science goals for the LUXE experiment planned at DESY [33, 34] and the E320 experiment at SLAC [35].
@@
In this asymmetric collision, a scattering event typically removes one x-ray and one high energy $\gamma$ photon from the beam and replaces it with two lower energy $\gamma$ photons. If sufficient numbers of $\gamma$ photons were to scatter in the x-ray field, we would therefore be able to detect the effect in the photon energy spectrum. This effect is illustrated in Fig. 1 for the idealised case of a 500 $\pm$ 25  MeV $\gamma$ photon beam scattering in a very dense field of 5 keV photons with $n_{x}=10^{27}$ photons mm-3 and length 1 cm. If however, insufficient scattering events occur to produce a measurable change in the spectrum, this allows us to place a limit on the photon-photon scattering cross-section.
@@
Figure 1: Illustration of the effect on the photon energy spectrum of copious photon-photon scattering of a 500 $\pm$ 25  MeV $\gamma$ photon beam in a very dense field of 5 keV photons with $n_{x}=10^{27}$ photons mm-3 and length 1 cm. Scattering in such a high density field results in a broadening and downshift of the $\gamma$ photon energy.
@@ 2 Experimental set-up
One of the Gemini laser pulses was used to generate a dense x-ray field by rapidly heating a 100 nm germanium (Ge) foil. As this solid Ge foil is heated, it turns into a plasma, leading to the emission of intense x-ray radiation predominantly due to M-L band transitions [54, 55].
The heating laser pulse had a duration of 40 ps (fwhm intensity) and a total energy of $10.7\pm 0.3$ J. It was focused to an elliptical spot using a distributive phase plate, with major and minor axes of $(217\pm 6)$ µm and $(77\pm 6)$ µm respectively, which contained 72% of the total energy. The Ge targets were mounted on a Kapton (C22H10N2O5) tape with a lower average atomic number, limiting the mass of Ge close to the interaction which is a potential noise source. A motorised tape-drive was used to change targets between shots.
To diagnose the x-ray field, a pinhole imaging system and crystal spectrometer were used. The pinhole imaging system gave an on-shot measure of both the emission spot size and the target alignment. The spectrometer used a flat, thallium acid phthalate (TlAP) crystal, with a spectral window of $\approx 700$ eV, centred at approximately $1.6$ keV (although the signal above approximately $1.5$ keV was supressed due to an aluminium filter). This spectral window is around the M-L band transitions of Ge. A measurement of the x-ray spectrum, averaged over 47 shots is shown in fig 3a).
@@ 2.2 $\gamma$ photon source
The $\gamma$ photon source was generated through bremsstrahlung emission, which first requires a beam of high energy electrons. To produce these electrons, one of the Gemini laser beams was focused into a 17.5 mm gas cell filled with helium and a 2% nitrogen dopant. The duration of the laser pulse was $45\pm 5$ fs (fwhm intensity) and the focal spot was $(44\pm 2)$ µm $\times$ $(53\pm 2)$ µm (fwhm intensity). The laser energy on target was $5.5\pm 0.6$ J, corresponding to a normalised vector potential $a_{0}=1.1\pm 0.2$. Through the laser wakefield acceleration mechanism [36], a beam of high energy electrons (energy up to $\approx 800$ MeV, charge $\approx 50$ pC ) was emitted from the gas cell.
These electrons then passed through a 0.5 mm thick bismuth (Bi) foil, acting as a bremsstrahlung converter. This emits a beam of high energy $\gamma$ photons with a similar duration to that of the driving laser pulse (i.e $\sim 50$ fs).
A calculation of the $\gamma$ photon spectrum produced in the experiment, based on Geant4 calculations of the bremsstrahlung conversion process for the measured electron beam spectrum is shown in figure 3b).
2.3 Spatio-temporal alignment
To achieve successful collisions between the two photon sources, it is necessary to overlap the sources in both space and time with sufficient precision. One of the advantages of this experiment is that the precision required to achieve collisions is not difficult to achieve. As the photon sources are driven by optical lasers it is straightforward to overlap the two sources with 1 ps and 10 µm precision by overlapping the drive laser pulses using fast diodes and optical imaging. Since the x-ray field size is $\sim 1\,$mm and the x-ray field duration is $\approx 40\,$ps, this is more than sufficient to achieve spatial and temporal overlap of the photon sources at the desired collision location.
2.4 Background
@@
The bremsstrahlung process used to generate the high-energy $\gamma$ rays also generates a large number of low energy, divergent $\gamma$ photons, which are a potential noise source in this experiment. If these divergent $\gamma$ photons were to interact with the Ge foil, or another part of the experimental setup, they would produce background through the Compton scattering process. To prevent this a 100 mm block of tungsten (W) with a 2 mm diameter hole drilled through the centre was used to collimate the $\gamma$ photon beam. This collimator effectively removes any $\gamma$ photon traveling at an angle greater than 10 mrad to the beam propagation direction.
@@
To further reduce the background, a 50 mm tungsten block was placed just off-axis, shadowing the Ge foil from the $\gamma$ photon beam. Placing such high Z-material close the $\gamma$ photon beam axis will itself generate a large number of background Bethe–Heitler pairs which would pass through the x-ray field and create a background signal through Compton scattering. Therefore, a 30 cm dipole magnet with a field strength of $B=1$ T, was used to remove these before the photon-photon interaction zone.
The target design and experimental geometry ensure that no plasma is in the path of the $\gamma$ ray beam when the $\gamma$ rays pass near the x-ray generating foil. This is achieved by mounting the germanium foil on a $4\,\mathrm{\mu m}$ kapton layer. The x-rays pass through the kapton to reach the collision point but the kapton layer prevents the expansion of any plasma towards the collision volume. Even without this kapton layer plasma expansion would not be an issue as the closest that the edge of the collimated $\gamma$ beam passes to the foil is 1 mm, 40 ps after the start of the laser pulse. Laser produced plasmas have ablation speeds of 100 – 10000 km s-1, so any plasma would only have expanded by 0.1 – 1 $\mu$m at the time of the arrival of the $\gamma$ photons.
Any effect of $\gamma$ photons interacting with material in the chamber, including the x-ray target foil are captured in the analysis by comparing the $\gamma$ spectrum on collision shots with null shots where the experimental geometry is identical but there is no laser incident on the x-ray target foil.
@@ 2.5 $\gamma$ photon spectrometer
The $\gamma$ photon spectrometer consisted of an array of $5\times 5$ mm crystals of CsI doped with thallium. The array was arranged in 47 columns ($z$) and 33 rows ($y$), with the long side of the crystals ($x$) oriented transversely to the propagation direction of the $\gamma$ rays ($z$). The deposition of energy by the incident $\gamma$-rays inside the array was captured by imaging the light emission by the crystals with an EMCCD camera. The deposition in the $z$-direction can be used to infer the $\gamma$-ray energy spectrum, while the response in the $y$-axis encodes information about the vertical divergence of the radiation. For our analysis we integrate the signal along the $y$-axis, treating each column of CsI crystals together. The spectrometer method is described in more detail in [31, 37].
The $\gamma$-photon spectrum is determined using a forward model based on a trial function. Bayesian inference is used to determine the best-fit parameters of the trial function and their uncertainty.
The detector was calibrated to remove systematic effects such as variations in light yield in each crystal due to crystal imperfections and misalignment, or effects produced by the imaging system. We compared the measured energy deposited in each crystal averaged over a large number of shots with that predicted in Geant4 using the average electron energy spectrum (as measured on a series of shots without the bismuth foil intercepting the electron beam) and generated a correction factor which can then be applied to each column of crystals.
@@
With the likelihood and priors set, we can use equation 3 to calculate the posterior. Given that the numerator involves a three-dimensional integral, it is most efficiently solved using a Markov chain Monte Carlo (MCMC) method. In figure 3c) we can see an example of this calculation performed on a randomly selected shot from the data set.
Figure 3: a) Measured x-ray photon spectrum, showing average and standard deviation measured over 47 shots, b) $\gamma$ photon spectrum (Geant4 simulations based on measured electron beam, showing average and standard deviation measured over 10 shots) c) Measured signal recorded by the CsI array from a randomly selected null shot (red circles). The signal is the intensity of light emitted by each column of CsI crystals (i.e eqn 1), normalised to the peak signal. The signal generated by the forward model of the detector for the average (black) and 100 randomly selected samples (grey) samples from the posterior distribution of $E_{c}$ is also shown. The mean value corresponds to $E_{c}=83.2\pm 3.2$ MeV.
@@ 3.1 $\gamma$ photon spectrum
We compare the distribution of $E_{c}$ on collision and null shots in various ways. Figure 4a) shows the histogram of the inferred value for the $\gamma$ photon spectrum $E_{c}$. The relatively small number of shots in each distribution means it is not immediately clear if differences between them are significant.
Figure 4b) shows the empirical cumulative distribution function (ecdf) of $E_{c}$ for the data. Also shown are 50 ecdfs for bootstrap samples of the data, these effectively represent the uncertainty in the ecdfs. The overlap between these ecdfs illustrates that there is no significant difference between the distributions. Figures 4c) and d) show the distribution of mean and standard deviation calculated from 100,000 bootstrap samples of the null and collision shot data. The fact that the distributions overlap further illustrates that there is no significant difference between the distribution of $E_{c}$ on null and collision shots.
Figure 4: a) Distribution of inferred value for the $\gamma$ photon spectrum $E_{c}$ for 32 null shots (blue) and 22 collision shots (red) b) Bootstrap estimate of the mean of $E_{c}$ for null shots (blue) and collision shots (red) c) Bootstrap estimate of the standard deviation of $E_{c}$ for null shots (blue) and collision shots (red)
@@
where $F_{\mathrm{N}}\left(E_{c}\right)$ and $F_{\mathrm{C}}\left(E_{c}\right)$ are the cumulative distribution functions for the null and collision shots respectively. The null hypothesis is accepted at the 95% confidence level if $D<0.378$[56] for data sets with 32 and 22 samples. The value obtained for our data set is D = 0.216, so we cannot reject the null hypothesis.
We can also calculate the two-sample Kolmogorov-Smirnov test statistic for a large number of bootstrap samples from the data to estimate the uncertainty in the KS test statistic (shown in Fig. 5). We find that the bulk of the distribution ($\approx 90\%$) lies below the critical value, providing further strong evidence that we cannot reject the null hypothesis, i.e. we must assume that collisions between $\gamma$ photons and the dense x-ray field did not produce a detectable difference in the energy spectrum of the $\gamma$ photons.
Figure 5: a) The empirical cumulative distribution (ecdf) function of $E_{c}$ for null shots (blue) and collision shots (red). The thick lines show the ecdf for the experimental data, the thin lines show ecdfs calculated from 50 bootstrap samples of the data. b) Distribution of the two sample Kolmogorov-Smirnov test statistic comparing the distribution of $E_{c}$ for null shots and collision shots from 100,000 bootstrap samples from the data. The red line shows the value above which the null hypothesis (that the null and collision shots are from the same distribution) can be rejected at the $\alpha=0.05$ significance level.
@@ 4 Conclusions
The data that support the findings of this study are openly available at the following URL/DOI: https://dx.doi.org/10.5281/zenodo.13767363.